At a certain college, there were 700 science majors, 200 engineering majros, and 800 business majors. If one student was selected at random, the probability that they are an engineering major is
Is it 200/1700 = 2/17?
A bookcase contains 3 statistics books and 5 biology books. If 2 books are chosen at random, the chance that both are statistic books is
I did: (3/8)(2/7) = 6/56 = 3/28
Why is it 2/7?
A box contains 24 transistors 4 of which are defective. If 4 transistors are randomly selected, what is the probability of selecting at least 1 defective transistor?
I did 20C4 / 24C4 = 1615/3542
3542/3542 - 1615/3542 = 1927/3542
A group of 10 male and 8 female students are planning to go out for pizza. If 50% of the male students go & 25% of the female students go, find the probability that a random student who goes out for pizza is female?
It has been reported that 60% of all cars on the highway are traveling at speeds of 60-70 mph. If the speed of three random cars are measured via radar, what is the probability that at least 1 car is driving at 60-70 mph?
A coin is tossed 5 times. Find the probability that the 6th toss is a tail, given that the first 5 tosses were all tails?
It's 1/2, but how do you show work for this?
In a second grade class containing 14 girls and 11 boys, 2 students are selected at random to give out math papers. What is the probability that the second student chose is a boy, given that the first one was a girl?

Answer by richard1234(5390)

1st question is correct.


2nd question -- 3/28 is correct. The 2/7 is in there because after one book is chosen, you have two biology books out of seven total. You could also do (3C2)/(8C2) = 3/28.

3rd question is correct.

4th question -- If I told you "5 males and 2 females go out for pizza" can you figure out the probability?

5th question -- Use the same method as in the third question, 1 minus the probability of none of them are driving 60-70 mph.

6th question -- Coin tosses are independent, so P(tail) = 1/2.

7th question -- Now we have 13 girls and 11 boys, what is P(boy)?